OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9

OpenStudy (lieutenantgeneral):

Reorder the terms: y = -9 + 2x Solving y = -9 + 2x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Simplifying y = -9 + 2x -2x+y=-9 2x-y=9 Standard form is when all the variables are on the same side and there are no fractions, so in this case, we move the 2x y=2x - 9 Subtract 2x from both sides The final answer is -2x + y = -9